# Relational Prompt-based Pre-trained Language Models for Social Event Detection

### v1

@@ 5.1.2. Online Evaluation
FinEvent $\mathrm{RPLM}_{SED}$ Improve
NMI
.$32\pm.02$
.$28\pm.00$
.$75\pm.02$
.$78\pm.02$
.$81\pm.02$
.$\underline{.82\pm.01}$
.$\textbf{.93}\pm\textbf{.01}.$
.$\uparrow.11$
AMI
.$32\pm.02$
.$03\pm.00$
.$74\pm.01$
.77±.01plus-or-minus.77.01.$7$01
.$81\pm.02$
.$\underline{.82\pm.01}$
.$\textbf{.93}\pm\textbf{.01}.$
.$\uparrow.11$
ARI
.$31\pm.02$
.$03\pm.00$
.$75\pm.02$
.$79\pm.02$
.$84\pm.02$
.$\underline{.86\pm.01}$
.$\textbf{.95}\pm\textbf{.01}.$
.$\uparrow.09$
5.2. Performance Evaluation on Low-Resource Language
Low-resource social event detection often faces the challenge of scarce training resources. Recent research introduced the CLKD (Ren et al., 2021) method, utilizing a knowledge distillation-based framework (Gupta et al., 2016) to address this challenge. Specifically, CLKD initially employs pre-trained word embedding algorithms (Mikolov et al., 2013; Pennington et al., 2014; Bojanowski et al., 2017) to generate initial embeddings of messages, followed by mapping the data of low-resource languages into the word embedding space of high-resource languages through linear (Lample et al., 2018; Artetxe et al., 2018) or non-linear (Mohiuddin et al., 2020) mapping techniques. Subsequently, a teacher model trained in a high-resource language (e.g., English) guides the student model in low-resource languages (e.g., Arabic) for learning and training. However, similar to other GNN-based methods, the backbone encoder of CLKD is still GNN. GNNs learn representations from explicit message graphs, within which some noisy edges are inevitably present. This results in the knowledge learned by the teacher model not being entirely effective and accurate, which may adversely affect the learning process of the student model. To validate our hypothesis, we conduct experiments on the low-resource Arabic-Twitter dataset. We choose TwHIN-BERT (TwHIN-BERT-large) as the backbone architecture for $\mathrm{RPLM}_{SED}$ and implement CLKD by using open-source code [^5] https://github.com/RingBDStack/CLKD. For other baseline methods, we use pre-trained word embedding algorithm [^6] https://github.com/bakrianoo/aravec to convert Arabic messages into initial embeddings before training. To ensure a fair comparison, all methods utilize the K-Means clustering method. As shown in Table 10, $\mathrm{RPLM}_{SED}$ outperforms the best results in the baseline by 13.41%, 13.41%, and 10.47% on NMI, AMI, and ARI, respectively. This is attributed to the inherent model advantage of $\mathrm{RPLM}_{SED}$ based on Multilingual Pretrained Language Models (MPLMs). Compared to GNNs, MPLMs can understand and encode different languages. This advantage enables $\mathrm{RPLM}_{SED}$ based on MPLMs to obtain better message embeddings when dealing with low-resource data, as well as stronger capabilities for encoding contextual messages. Secondly, the pairwise learning mechanism, by consistently encoding both the content and structural information of messages, leverages more effective information. This enables the model to obtain robust and stable message representations, even in situations where training resources are limited.
Figure 4. Cluster visualization of message representations in the detection stage.
@@ 5.2. Performance Evaluation on Low-Resource Language
Dataset Metrics
PPGCN
KPGNN
QSGNN
FinEvent
MVGAN
ETGNN
$\mathrm{UCL}_{SED}$
$\mathrm{RPLM}_{SED}$
Improve
Events 2012_100 ACC
.$64\pm.00$
.$73\pm.01$
.$75\pm.01$
.$81\pm.01$
.$82\pm.00$
.$87\pm.01$
.$92\pm.01$
.$\textbf{.96}\pm\textbf{.01}.$
.$\uparrow.04$
F1
.55±.00plus-or-minus.55.00.$5$00
.$60\pm.00$
.$69\pm.01$
.$80\pm.01$
.$82\pm.01$
.$87\pm.00$
.$92\pm.01$
.$\textbf{.97}\pm\textbf{.01}.$
.$\uparrow.05$
Events 2018_100 ACC
.$71\pm.01$
.77±.01plus-or-minus.77.01.$7$01
.77±.01plus-or-minus.77.01.$7$01
.$63\pm.00$
.$68\pm.01$
.$61\pm.01$
.$79\pm.01$
.$\textbf{.90}\pm\textbf{.01}.$
.$\uparrow.11$
F1
.$51\pm.00$
.$62\pm.01$
.$75\pm.01$
.$63\pm.01$
.$68\pm.01$
.$60\pm.01$
.$79\pm.01$
.$\textbf{.92}\pm\textbf{.01}.$
.$\uparrow.12$
5.3. Performance Evaluation on Long-tail Recognition Task
In real-world social data, the characteristics of a long-tail distribution are particularly evident. There are few head-dominated event classes, whereas low-frequency tail classes are more numerous. This imbalance in training data makes it difficult for tail events to be correctly classified. In recent research, ETGNN (Ren et al., 2022b) employs the Dempster-Shafer theory (Sentz and Ferson, 2002) to integrate the uncertainty and temporal information of graphs constructed from different perspectives, thereby enhancing the accuracy of event detection. Furthermore, building upon ETGNN, the $\mathrm{UCL}_{SED}$ (Ren et al., 2023) framework is proposed. Compared to ETGNN, $\mathrm{UCL}_{SED}$ introduces uncertainty-guided contrastive learning loss, pushing boundary-ambiguous event classes further away to improve the distinguishability of representations. We also compare with MVGAN, which learns message representations from semantic and temporal views of messages and proposes a hashtag-based multi-view graph attention mechanism to fuse representations. We conduct experiments on two long-tail datasets, Event2012_100 and Event2018_100. It is noteworthy that to accommodate the classification task, an event classifier is additionally incorporated into $\mathrm{RPLM}_{SED}$. For the acquisition of message labels, we first filter out unreliable labels from their corresponding similarity sets based on similarity and then determine their final labels through a voting mechanism. In terms of evaluation metrics, we select Accuracy (ACC) and F1 value (F1), which are widely used in classification tasks. As illustrated in Table 11, compared with the optimal results among the baseline methods, $\mathrm{RPLM}_{SED}$ exhibits an improvement of 4.34% in ACC and 5.43% in F1, respectively, on Event2012_100. On Event2018_100, the improvements are 19.92% and 16.46%, respectively. Recalling the pairwise message modeling process, we sample multiple messages for each message to construct message pairs. Utilizing the multi-relational prompt-based pairwise message learning mechanism, each message is learned multiple times from different message pairs. This allows messages of tail events to effectively learn a common representation of similar messages in terms of semantic and structural relations, as well as the subtle differences between different classes of messages. Additionally, the clustering constraint forces the representations of each event class to be as dispersed as possible in the latent space while ensuring that representations of messages within the same class are as close as possible. This further enhances the distinguishability of message representations, thereby enabling the event classifier to predict labels with greater accuracy. Furthermore, for each message, by filtering out unreliable labels from its corresponding set of candidate labels based on similarity, the label subsequently obtained through voting from the remaining labels is more reliable.
Figure 5. Overall Performance of Various PLM-based Models on Events2012 in Online Scenario.
@@ 5.4. PLM-based Model for Social Event Detection
In the preceding sections, we comprehensively compare $\mathrm{RPLM}_{SED}$ against the currently dominant social event detection methods based on GNNs across various scenarios. To our knowledge, $\mathrm{RPLM}_{SED}$ is the inaugural model leveraging PLMs to achieve high-quality social event detection in diverse scenarios, including offline, online, low-resource settings, and long-tail distributions. To further explore the performance of Pre-trained Language Models (PLMs) in the social event detection tasks, we design experiments utilizing the pre-trained word embedding technology Word2Vec and two mainstream PLMs, BERT and RoBERTa. Specifically, for Word2Vec, consistent with previous methods, we conduct clustering based on the average embeddings of message content. Additionally, for the original BERT and RoBERTa, we implement two distinct strategies: without fine-tuning and fine-tuning. Furthermore, we propose variant models BERT-LR and RoBERTa-LR for $\mathrm{RPLM}_{SED}$ based on BERT and RoBERTa, respectively, which incorporate fine-tuning through LoRA. In practical applications, given the infeasibility of the K-Means clustering method due to its requirement for pre-specifying the number of events, we consistently employed the HDBSCAN clustering method. It is important to note that $\mathrm{RPLM}_{SED}$ is predicated on RoBERTa-base, and similarly, all other models also utilize their respective base versions.
@@
As illustrated in Figure 5, among the methods for directly obtaining message representations from Word2Vec, BERT, and RoBERTa, BERT demonstrates the relatively best performance. This may be attributed to the higher compatibility between the knowledge acquired during the BERT pre-training process and the Event2012 dataset in the absence of model fine-tuning. Upon fine-tuning, BERT-FT and RoBERTa-FT exhibit significant enhancements, achieving nearly equivalent model performances. Nonetheless, a discernible performance gap remains when compared to $\mathrm{RPLM}_{SED}$, attributable to their exclusive reliance on the semantic information of messages while neglecting the abundant structural information within the messages. Moreover, we observe that the performances of the variants BERT-LR and RoBERTa-LR fall short of those achieved through global fine-tuning of the models. In online event detection scenarios, models necessitate continual updates to incorporate new knowledge while concurrently sustaining an existing knowledge base. Utilizing LoRA to fine-tune models can effectively decrease the number of parameters requiring adjustments during the fine-tuning process, which aids in enhancing the efficiency of model updates and reducing computational costs. However, this strategy of reducing the volume of trainable parameters may adversely affect the model’s long-term capabilities for knowledge learning, retention, and expansion, consequently leading to a decline in model performance. Overall, RPLM’s model performance significantly surpasses that of other methods, further illustrating the efficacy of the multi-relational prompt-based pairwise message learning mechanism. On the other hand, the experiments also validate the feasibility of utilizing PLMs for social event detection tasks.
@@ 5.5. $\mathrm{RPLM}_{SED}$ with Different PLMs
Figure 6. $\mathrm{RPLM}_{SED}$ performance comparison based on different PLMs.
As depicted in Figure 6, compared to $\mathrm{RPLM}_{SED}$ based on RoBERTa-base, the model based on RoBERTa-large exhibits overall improvements of $3.23\%$, $3.45\%$, and $4.68\%$ on the NMI, AMI, and ARI, respectively. This is because, compared to RoBERTa-base, RoBERTa-large possesses a larger model size, more trainable parameters, and a broader training dataset, resulting in superior model performance. Naturally, the $\mathrm{RPLM}_{SED}$ based on RoBERTa-large requires greater computational resources and training time. The primary distinction between Twitter-RoBERTa-large and RoBERTa-large lies in the difference in training corpora. Twitter-RoBERTa-large is specifically trained on Twitter data, whereas RoBERTa-large undergoes pre-training on a broader spectrum of text data. These two models have the same model scale and have been further fine-tuned on the same dataset, hence their overall performance is essentially consistent. When contrasting RoBERTa-base with BERT-base-uncased, two distinct types of pre-trained language models, we observe that $\mathrm{RPLM}_{SED}$ based on RoBERTa-base exhibits superior performance. In terms of overall performance, the $\mathrm{RPLM}_{SED}$ based on RoBERTa-base surpasses the $\mathrm{RPLM}_{SED}$ based on BERT-base-uncased by 4.36%, 4.81%, and 12.15% in NMI, AMI, and ARI, respectively. This advantage can be attributed to the pre-training of the RoBERTa-base on a larger dataset and optimized training strategies. Notably, even the $\mathrm{RPLM}_{SED}$ based on BERT-base-uncased outperforms the strongest baseline model, FinEvent, leading by 2.81% in NMI, 2.51% in AMI, and 8.75% in ARI.
@@ 5.6. Ablation Study
Similarity-based Representation Aggregation. We investigate the variations in model performance under different representation aggregation methods, including similarity-based representation aggregation method, top-k sampling representation aggregation method, and all representation average aggregation methods. For a set of candidate feature representations obtained for each message, under the top-k sampling representation aggregation method, we only sample the top 20 candidate representations with the highest similarities for averaging. This approach has certain limitations. A smaller value of $k$ may lead to message representations that are not comprehensive enough. Conversely, a larger $k$ may include candidate representations extracted from negative sample pairs, which often carry noise and irrelevant information, thereby affecting the quality of the final message representations. For the all-representation average aggregation method, averaging all candidate representations of messages may lead to overly smoothed representations and is also susceptible to noise from candidate representations extracted from negative sample pairs. Consequently, among the three methods, it performs the worst. In contrast, the similarity-based representation aggregation approach averages candidate representations of messages with high confidence, effectively filtering out noisy candidate representations extracted from negative sample pairs while preventing message representations from becoming overly smoothed. Overall, the similarity-based representation aggregation approach outperforms the top-k sampling representation aggregation by $1.04\%$, $1.05\%$, and $2.89\%$ in NMI, AMI, and ARI, respectively. Additionally, it surpassed the all representation average aggregation method by 3.68%, 3.84%, and 8.16% on NMI, AMI, and ARI, respectively.

### v2

@@ 5.2. Performance Evaluation on Low-Resource Language
Metrics PP-GCN EventX KPGNN QSGNN CLKD FinEvent $\mathrm{RPLM}_{SED}$ Improve
NMI .$32\pm.02$$28\pm.00$$75\pm.02$$78\pm.02$$81\pm.02$$\underline{.82\pm.01}$$\textbf{.93}\pm\textbf{.01}.$$\uparrow.11$
AMI .$32\pm.02$$03\pm.00$$74\pm.01$$7$01 .$81\pm.02$$\underline{.82\pm.01}$$\textbf{.93}\pm\textbf{.01}.$$\uparrow.11$
ARI .$31\pm.02$$03\pm.00$$75\pm.02$$79\pm.02$$84\pm.02$$\underline{.86\pm.01}$$\textbf{.95}\pm\textbf{.01}.$$\uparrow.09$
For other baseline methods, we use pre-trained word embedding algorithm [^7] https://github.com/bakrianoo/aravec to convert Arabic messages into initial embeddings before training. To ensure a fair comparison, all methods utilize the K-Means clustering method.
As shown in Table 10, $\mathrm{RPLM}_{SED}$ outperforms the best results in the baseline by 13.41%, 13.41%, and 10.47% on NMI, AMI, and ARI, respectively. This is attributed to the inherent model advantage of $\mathrm{RPLM}_{SED}$ based on Multilingual Pretrained Language Models (MPLMs). Compared to GNNs, MPLMs can understand and encode different languages. This advantage enables $\mathrm{RPLM}_{SED}$ based on MPLMs to obtain better message embeddings when dealing with low-resource data, as well as stronger capabilities for encoding contextual messages. Secondly, the pairwise learning mechanism, by consistently encoding both the content and structural information of messages, leverages more effective information. This enables the model to obtain robust and stable message representations, even in situations where training resources are limited.
5.3. Performance Evaluation on Long-tail Recognition Task
In real-world social data, the characteristics of a long-tail distribution are particularly evident. There are few head-dominated event classes, whereas low-frequency tail classes are more numerous. This imbalance in training data makes it difficult for tail events to be correctly classified. In recent research, ETGNN (Ren et al., 2022b) employs the Dempster-Shafer theory (Sentz and Ferson, 2002) to integrate the uncertainty and temporal information of graphs constructed from different perspectives, thereby enhancing the accuracy of event detection. Furthermore, building upon ETGNN, the $\mathrm{UCL}_{SED}$ (Ren et al., 2023) framework is proposed. Compared to ETGNN, $\mathrm{UCL}_{SED}$ introduces uncertainty-guided contrastive learning loss, pushing boundary-ambiguous event classes further away to improve the distinguishability of representations. We also compare with MVGAN (Cui et al., 2021), which learns message representations from semantic and temporal views of messages and proposes a hashtag-based multi-view graph attention mechanism to fuse representations.
We conduct experiments on two long-tail datasets, Events2012_100 and Events2018_100. It is noteworthy that to accommodate the classification task, an event classifier is additionally incorporated into $\mathrm{RPLM}_{SED}$. For the acquisition of message labels, we first filter out unreliable labels from their corresponding similarity sets based on similarity and then determine their final labels through a voting mechanism. In terms of evaluation metrics, we select Accuracy (ACC) and F1 value (F1), which are widely used in classification tasks. As illustrated in Table 11, compared with the optimal results among the baseline methods, $\mathrm{RPLM}_{SED}$ exhibits an improvement of 4.34% in ACC and 5.43% in F1, respectively, on Events2012_100. On Events2018_100, the improvements are 19.92% and 16.46%, respectively. Recalling the pairwise message modeling process, we sample multiple messages for each message to construct message pairs. Utilizing the multi-relational prompt-based pairwise message learning mechanism, each message is learned multiple times from different message pairs. This allows messages of tail events to effectively learn a common representation of similar messages in terms of semantic and structural relations, as well as the subtle differences between different classes of messages. Additionally, the clustering constraint forces the representations of each event class to be as dispersed as possible in the latent space while ensuring that representations of messages within the same class are as close as possible. This further enhances the distinguishability of message representations, thereby enabling the event classifier to predict labels with greater accuracy. Furthermore, for each message, by filtering out unreliable labels from its corresponding set of candidate labels based on similarity, the label subsequently obtained through voting from the remaining labels is more reliable.
@@ 5.3. Performance Evaluation on Long-tail Recognition Task
Dataset Metrics PPGCN KPGNN QSGNN FinEvent MVGAN ETGNN $\mathrm{UCL}_{SED}$ $\mathrm{RPLM}_{SED}$ Improve
Events 2012_100 ACC .$64\pm.00$$73\pm.01$$75\pm.01$$81\pm.01$$82\pm.00$$87\pm.01$$92\pm.01$$\textbf{.96}\pm\textbf{.01}.$$\uparrow.04$
F1 .55±.00plus-or-minus.55.00.$5$00 .$60\pm.00$$69\pm.01$$80\pm.01$$82\pm.01$$87\pm.00$$92\pm.01$$\textbf{.97}\pm\textbf{.01}.$$\uparrow.05$
Events 2018_100 ACC .$71\pm.01$$7$01 .77±.01plus-or-minus.77.01.$7$01 .$63\pm.00$$68\pm.01$$61\pm.01$$79\pm.01$$\textbf{.90}\pm\textbf{.01}.$$\uparrow.11$
F1 .$51\pm.00$$62\pm.01$$75\pm.01$$63\pm.01$$68\pm.01$$60\pm.01$$79\pm.01$$\textbf{.92}\pm\textbf{.01}.$$\uparrow.12$
@@ 5.4. PLM-based Model for Social Event Detection
In the preceding sections, we comprehensively compare $\mathrm{RPLM}_{SED}$ against the currently dominant social event detection methods based on GNNs across various scenarios. To our knowledge, $\mathrm{RPLM}_{SED}$ is the inaugural model leveraging PLMs to achieve high-quality social event detection in diverse scenarios, including offline, online, low-resource settings, and long-tail distributions. To further explore the performance of Pre-trained Language Models (PLMs) in the social event detection tasks, we design experiments utilizing the pre-trained word embedding technology Word2Vec and two mainstream PLMs, BERT and RoBERTa. Specifically, for Word2Vec, BERT, and RoBERTa, which only use the message content as input, the final representation is obtained by averaging the embeddings of each word in the message. Additionally, we implement two distinct strategies for the original BERT and RoBERTa without fine-tuning and fine-tuning strategies. Furthermore, we propose variant models BERT-LR and RoBERTa-LR for $\mathrm{RPLM}_{SED}$ based on BERT and RoBERTa, respectively, which incorporate fine-tuning through LoRA. In practical applications, given the infeasibility of the K-Means clustering method due to its requirement for pre-specifying the number of events, we consistently employed the HDBSCAN clustering method. It is important to note that $\mathrm{RPLM}_{SED}$ is predicated on RoBERTa-base, and similarly, all other models also utilize their respective base versions.
@@
As illustrated in Figure 15, among the methods for directly obtaining message representations from Word2Vec, BERT, and RoBERTa, BERT demonstrates the relatively best performance. This may be attributed to the higher compatibility between the knowledge acquired during the BERT pre-training process and the Events2012 dataset in the absence of model fine-tuning. Upon fine-tuning, BERT-FT and RoBERTa-FT exhibit significant enhancements, achieving nearly equivalent model performances. Nonetheless, a discernible performance gap remains when compared to $\mathrm{RPLM}_{SED}$, attributable to their exclusive reliance on the semantic information of messages while neglecting the abundant structural information within the messages. Moreover, we observe that the performances of the variants BERT-LR and RoBERTa-LR fall short of those achieved through global fine-tuning of the models. In online event detection scenarios, models necessitate continual updates to incorporate new knowledge while concurrently sustaining an existing knowledge base. Utilizing LoRA to fine-tune models can effectively decrease the number of parameters requiring adjustments during the fine-tuning process, which aids in enhancing the efficiency of model updates and reducing computational costs. However, this strategy of reducing the volume of trainable parameters may adversely affect the model’s long-term capabilities for knowledge learning, retention, and expansion, consequently leading to a decline in model performance. Overall, RPLM’s model performance significantly surpasses that of other methods, further illustrating the efficacy of the multi-relational prompt-based pairwise message learning mechanism. On the other hand, the experiments also validate the feasibility of utilizing PLMs for social event detection tasks.
Figure 15. Overall Performance of Various PLM-based Models on Events2012 in Online Scenario. The horizontal axis represents the average metric values of different methods, and the best results are bolded.
@@ 5.5. $\mathrm{RPLM}_{SED}$ with Different PLMs
As depicted in Figure LABEL:fig:diff-plm, compared to $\mathrm{RPLM}_{SED}$ based on RoBERTa-base, the model based on RoBERTa-large exhibits overall improvements of $3.23\%$, $3.45\%$, and $4.68\%$ on the NMI, AMI, and ARI, respectively. Compared to RoBERTa-base, RoBERTa-large possesses a larger model size, more trainable parameters, and a broader training dataset, resulting in superior model performance. Naturally, the $\mathrm{RPLM}_{SED}$ based on RoBERTa-large requires greater computational resources and training time. The primary distinction between Twitter-RoBERTa-large and RoBERTa-large lies in the difference in training corpora. Twitter-RoBERTa-large is specifically trained on Twitter data, whereas RoBERTa-large undergoes pre-training on a broader spectrum of text data. These two models have the same model scale and have been further fine-tuned on the same dataset, hence their overall performance is essentially consistent. When contrasting RoBERTa-base with BERT-base-uncased, two distinct types of pre-trained language models, we observe that $\mathrm{RPLM}_{SED}$ based on RoBERTa-base exhibits superior performance. In terms of overall performance, the $\mathrm{RPLM}_{SED}$ based on RoBERTa-base surpasses the $\mathrm{RPLM}_{SED}$ based on BERT-base-uncased by 4.36%, 4.81%, and 12.15% in NMI, AMI, and ARI, respectively. This advantage can be attributed to the pre-training of the RoBERTa-base on a larger dataset and optimized training strategies. Notably, even the $\mathrm{RPLM}_{SED}$ based on BERT-base-uncased outperforms the strongest baseline model, FinEvent, leading by 2.81% in NMI, 2.51% in AMI, and 8.75% in ARI.
5.6. Time Consumption of $\mathrm{RPLM}_{SED}$ based on Different Scale PLMs
This section reports the time consumption of different parts of $\mathrm{RPLM}_{SED}$ based on RoBERTa-large and RoBERTa-base on the Events2012. Figure 19 shows the message number of each block, the time consumption for pairwise messages modeling during the detection phase, and the execution time for detection operations. $M_{0}$ is solely used for initial model training and incurs no time consumption during detection. Additionally, Table 12 presents statistics on message numbers within each window, the time required for pairwise message modeling during training, and the time consumption for model training or maintenance in each window. Both Figures 19 and Table 12 show that the time expenditure for model training or detection correlates linearly with the number of messages in message blocks. Secondly, whether in the training or detection phase, the time spent on pairwise message modeling is significantly less than the time spent on model training or detection. Furthermore, the training and detection time of $\mathrm{RPLM}_{SED}$ based on RoBERTa-large is approximately 3.2 times that of $\mathrm{RPLM}_{SED}$ based on RoBERTa-base. This is attributed to RoBERTa-large’s larger hidden dimensions and more Transformer layers, which enhance its performance compared to RoBERTa-base-based $\mathrm{RPLM}_{SED}$. It should be noted that in addition to the employed PLM scale, factors such as message sampling numbers and hardware environment also markedly influence the overall time consumption of the model.
Figure 19. The number of messages in each message block (histogram) and the time consumption of pairwise modeling and detection for each block (line chart) in Events2012.
Table 12. The number of messages in each window and the time consumption of pairwise modeling and training (or maintenance) for each window in Events2012.
Message Number
Window $M_{0}$ $M_{1}-M_{3}$ $M_{4}-M_{6}$ $M_{7}-M_{9}$ $M_{10}-M_{12}$ $M_{13}-M_{15}$ $M_{16}-M_{18}$ $M_{19}-M_{21}$
Number 20254 12048 5120 8201 5565 7477 5473 4702
Time Consumption (min)
Window $M_{0}$ $M_{1}-M_{3}$ $M_{4}-M_{6}$ $M_{7}-M_{9}$ $M_{10}-M_{12}$ $M_{13}-M_{15}$ $M_{16}-M_{18}$ $M_{19}-M_{21}$
Pairwise Modeling 4.55 3.53 0.52 1.52 0.57 1.20 0.79 0.47
RoBERTa-base 55.06 32.88 14.73 22.28 15.08 19.95 14.7 12.58
RoBERTa-large 175.20 108.25 46.35 72.24 50.78 65.23 48.80 40.45
@@ 5.7. Ablation Study
Similarity-based Representation Aggregation. We investigate the variations in model performance under different representation aggregation methods, including similarity-based representation aggregation method, top-k sampling representation aggregation method, and all representation average aggregation method. For a set of candidate feature representations obtained for each message, under the top-k sampling representation aggregation method, we only sample the top 20 candidate representations with the highest similarities for averaging. This approach has certain limitations. A smaller value of $k$ may lead to message representations that are not comprehensive enough. Conversely, a larger $k$ may include candidate representations extracted from negative sample pairs, which often carry noise and irrelevant information, thereby affecting the quality of the final message representations. For the all-representation average aggregation method, averaging all candidate representations of messages may lead to overly smoothed representations and is also susceptible to noise from candidate representations extracted from negative sample pairs. Consequently, among the three methods, it performs the worst. In contrast, the similarity-based representation aggregation approach averages candidate representations of messages with high confidence, effectively filtering out noisy candidate representations extracted from negative sample pairs while preventing message representations from becoming overly smoothed. Overall, the similarity-based representation aggregation approach outperforms the top-k sampling representation aggregation by $1.04\%$, $1.05\%$, and $2.89\%$ in NMI, AMI, and ARI, respectively. Additionally, it surpassed the all representation average aggregation method by 3.68%, 3.84%, and 8.16% on NMI, AMI, and ARI, respectively.